A car travels at a constant speed of 15m/s. How many miles does the car travel in one hour?
Answer by mananth(16946)

15m/s =
1m=3.281 ft
15m=15*3.281 ft=49.2 ft
49.2ft = 49.2/5280=0.0093 miles
15m/s= 0.0093 mile/s
0.0093*3600= 33.54 miles/hour
m.ananth@hotmail.ca